On the improvement of the TDSCF method for a reaction coordinate coupled to a heat bath.

Billing G. D. Jolicard G.

Abstract

We have investigated the dynamics of a reaction coordinate coupled to a heat bath of harmonic oscillators and show that it is possible to introduce a simple correction to the SCF scheme such that good agreement with exact numbers is obtained for the tunneling probability.